OpenStudy (anonymous):

Perform the following operations. 1.) [(i)(i^2)(i^3)...(i^999)] / i + i^2 +i^3 + ... + i^999 2.) (...(((i)^1)^2)^3)...)^2014

OpenStudy (anonymous):

Is \(i\) the imaginary number \(i=\sqrt{-1}\) ?

OpenStudy (anonymous):

@SithsAndGiggles Yes.

OpenStudy (anonymous):

Okay, so for the first one, you have \[\large\frac{i\times i^2\times\cdots\times i^{999}}{i+i^2+\cdots+i^{999}}\] For the numerator, use the property that for any \(b>0\), you have \[\large b^a\times b^c=b^{a+c}\] which means the numerator is rewritten as \[\large i\times i^2\times\cdots\times i^{999}=i^{1+2+\cdots+999}\] The exponent is the sum of consecutive integers 1 to 999. Recall this useful formula for evaluating this sum: \[\large\sum_{k=1}^nk=1+2+\cdots+n=\frac{n(n+1)}{2}\] When \(n=999\), you have \[\large i^{1+\cdots+999}=i^{(999(999+1))/2}=i^{499500}\] Since 499500 is a multiple of 4, you can reduce this to \[\large i^{499500}=\left(i^4\right)^{124875}=1^{124875}=1\] So we've found that \[\large\frac{i\times i^2\times\cdots\times i^{999}}{i+i^2+\cdots+i^{999}}=\frac{1}{i+i^2+\cdots+i^{999}}\] Does this make sense so far?

OpenStudy (anonymous):

@SithsAndGiggles Yes. How do you solve the denominator?

OpenStudy (anonymous):

I'm back, sorry for the wait. I see two possible ways to rewrite the denominator. For one way, I think you can use the formula for finite geometric sums. \[\large \sum_{i=1}^nar^{i-1}=\frac{a(1-r^n)}{1-r}\] In this case, \(r=i\). The denominator is \[\large \begin{align*}i^{2-1}+i^{3-1}+\cdots+i^{1000-1}&=\sum_{k=2}^{1000}i^{k-1}\\&=\sum_{k=1}^{1000}i^{k-1}-1\\&=\frac{1-i^{1000}}{1-i}-1\\ &=\frac{1-\left(i^4\right)^{250}}{1-i}-1\\ &=\frac{1-1^{250}}{1-i}-1\\ &=0-1\\ &=-1 \end{align*}\] This means the first expression in your question reduces to \(\dfrac{1}{-1}=-1\). The other way would have been to count how many multiples of \(i\), \(i^2\), \(i^3\), and \(i^4\) there are, but I think that takes a bit longer.

OpenStudy (anonymous):

@SithsAndGiggles I'm sorry but I don't get the solution... I solved it earlier using the multiples and it was long, and I think your solution is shorter. How did you solve that?

OpenStudy (anonymous):

All the steps are listed in my previous comment. If you don't know the geometric sum formula, stick to the multiples method.

OpenStudy (anonymous):

@SithsAndGiggles Why did you put "-1" in the second step?

OpenStudy (anonymous):

That's because the formula for the summation applies to a starting index of \(k=1\), not \(k=2\). Note that \[\large\sum_{k=2}^{1000}i^{k-1}=i^1+i^2+\cdots+i^{999}\] whereas \[\large\sum_{\color{red}{k=1}}^{1000}i^{k-1}=\color{red}{i^0}+i^1+i^2+\cdots+i^{999}\] Since \(i^0=1\), you have \[\large\sum_{\color{red}{k=1}}^{1000}i^{k-1}=\color{red}{1}+i^1+i^2+\cdots+i^{999}\] and so we move this 1 to the left side: \[\large\begin{align*}\sum_{\color{red}{k=1}}^{1000}i^{k-1}-\color{red}1&=i^1+i^2+\cdots+i^{999}\\ &=\sum_{k=2}^{1000}i^{k-1} \end{align*}\]

OpenStudy (anonymous):

For the second question, consider the first four powers: \[\large \left(\left(\left(\left(\left(\left(i\right)^1\right)^2\right)^3\right)^4\right)^{\cdots}\right)^{2014}\] Recall the property of exponents: \[\large \left(a^b\right)^c=a^{bc}=a^{cb}=\left(a^c\right)^b\] This means we can rearrange the order of the exponents so that \(i\) is raised to the fourth power first: \[\large \left(\left(\left(\left(\left(\left(i\right)^\color{red}4\right)^2\right)^3\right)^\color{red}1\right)^{\cdots}\right)^{2014}\] Since \(i^4=1\), you have \[\large \left(\left(\left(\left(1^2\right)^3\right)^1\right)^{\cdots}\right)^{2014}=1\]